OpenStudy (anonymous):

Now they want x intercept.. is that the slope part. for instance.. If 2x + y = 7, then the x-intercept of the graph of this equation is: -2x?

OpenStudy (anonymous):

if you want the x intercept you just insert a zero for y.. so you get x = 7/2

OpenStudy (chaise):

Not quiet - At the x intercept the y point on your graph is 0 Change this to y=mx+b format and then make y = 0. :)

OpenStudy (anonymous):

okay let me work it

OpenStudy (chaise):

You don't even have to put it into point-slope formula. You can just make y = 0. When any line cuts the x axis the value for it is (x,0).

OpenStudy (anonymous):

2x + y = 7 ===> let y=0 ===> 2x=7 ===> x=7/2

OpenStudy (anonymous):

so 0 = -2x +7 -7 = -2x so x = 7/2?

OpenStudy (chaise):

Correct.

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

ok now we're cookin

OpenStudy (anonymous):

great